mathslover (mathslover):

The value of \[\lim_{x \rightarrow 0} \cfrac{\int_{0} ^{x^2} \sec^2 t dt }{x \sin x} = ? \]

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

First solve the integration part, the integration of sec^2 x dx is ?

mathslover (mathslover):

tanx + C (it is definite then remove C)

OpenStudy (vishweshshrimali5):

Good

OpenStudy (vishweshshrimali5):

Or you can use LH rule also

OpenStudy (anonymous):

Hopital will work also

mathslover (mathslover):

so if I use LH rule... then the numerator will become \(sec^2 t dt\) ?

OpenStudy (vishweshshrimali5):

No

OpenStudy (unklerhaukus):

Maclaurin

OpenStudy (vishweshshrimali5):

Remember this: *sigh* I don't know latex codes...

mathslover (mathslover):

Okay.. I thought integration is anti-derivative and so, if I differentiate the integral, it will give the function inside it.

ganeshie8 (ganeshie8):

chain rule - bounds matter !

ganeshie8 (ganeshie8):

\[\large \dfrac{d}{dx} \int \limits_{0}^{x^2} f(t) dt = f(x^2)\times (x^2)' \]

OpenStudy (vishweshshrimali5):

\[\cfrac{d}{dx}\int\limits_{f(x)}^{g(x)}\alpha(t).dt = \alpha(g(x)).g'(x) - \alpha(f(x)).f'(x)\]

OpenStudy (vishweshshrimali5):

Huhh I ultimately wrote it...

OpenStudy (vishweshshrimali5):

Well @ganeshie8 has put the values in the general one. You can check it @mathslover

mathslover (mathslover):

Yeah.. so I can write what @ganeshie8 wrote as : \(f(x^2) \times 2x\) And for the question : \(\sec^2 (x^2) \times 2x\) ... will become the numerator? is that right?

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (vishweshshrimali5):

You could have obtained the same thing, if you had first solved the integral and then applied the LH rule.

mathslover (mathslover):

SO, I get : \[ \lim_{x\rightarrow 0} \cfrac{\sec^2 (x^2) \times 2x}{ \cos x}\]

mathslover (mathslover):

Right?

mathslover (mathslover):

Oh no... sorry

mathslover (mathslover):

\(\sin x + x \cos x\) will be the denominator, right?

OpenStudy (vishweshshrimali5):

yeah

OpenStudy (vishweshshrimali5):

apply LH rule again

mathslover (mathslover):

Okay.. give me one minute :-)

OpenStudy (vishweshshrimali5):

take your time.

mathslover (mathslover):

Its a little bit complicated :/

ganeshie8 (ganeshie8):

\[\lim_{x \rightarrow 0} \cfrac{\int_{0} ^{x^2} \sec^2 t dt }{x \sin x} = \lim_{x \rightarrow 0} \cfrac{\tan(x^2) }{x \sin x} = \lim_{x \rightarrow 0} \cfrac{\dfrac{\tan (x^2)}{x^2} }{\dfrac{\sin x}{x}} \]

ganeshie8 (ganeshie8):

the first idea was simple ^ :)

mathslover (mathslover):

@ganeshie8 yes I got 1 from there already

OpenStudy (vishweshshrimali5):

yeah that's a better method.

OpenStudy (vishweshshrimali5):

Also, if you had obtained sin x + xcos x in the denominator then divide both numerator and denominator by x and apply this : lim x--> 0 sinx/x = 1

mathslover (mathslover):

2 sec^2 (t) tan (t)

mathslover (mathslover):

@vishweshshrimali5

mathslover (mathslover):

Yeah.. I got how you didn't use LH rule again. But, any way to go on from there? ^

OpenStudy (vishweshshrimali5):

you differentiated the numerator wrongly

mathslover (mathslover):

o.O Yeah..

OpenStudy (vishweshshrimali5):

remembe derivative sec (x) is sec(x)tan(x)

mathslover (mathslover):

I thought that there is sin x in the denominator :(

mathslover (mathslover):

Bhaiya, is the numerator right?

OpenStudy (vishweshshrimali5):

Nope

OpenStudy (vishweshshrimali5):

Neither numerator nor denominator are right

mathslover (mathslover):

Fine got it..

OpenStudy (vishweshshrimali5):

first note the result of first application of LH rule

mathslover (mathslover):

this is what I have..

OpenStudy (vishweshshrimali5):

I would advise you not to use this method. \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Also, if you had obtained sin x + xcos x in the denominator then divide both numerator and denominator by x and apply this : lim x--> 0 sinx/x = 1 \(\color{blue}{\text{End of Quote}}\) See the limit you obtained after applying the LH Rule for first time. Divide the numerator and denominator by x. And solve. You will get your answer

mathslover (mathslover):

Yep.. I understand that. But, I was just checking my differentiation skills

mathslover (mathslover):

the differentiation done by me there is right or wrong? can u please tell this?